OpenStudy (anonymous):

Solve the equation Three x minus six equals fifteen for x. x equals three x equals seven x equals twenty-seven x equals sixty-three

OpenStudy (anonymous):

State your words mathematically: 3x-6=15 Now add 6 to both sides of your equation.

OpenStudy (anonymous):

ok um.. well 6 + 3 is 9 and then 6 + 15 is 21

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

There is no 6+3 You can only add like terms. so you can add 10 to 20, and you can add 14y to 27y, but you can't add 10 to 27y So all you have by adding 6 to both sides is: 3x-6+6 = 15+6 So let's take it from here... :-)

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

So what do you get if you simplify: 3x -6 +6 = 15+6

OpenStudy (anonymous):

idk srry I really bad at math

OpenStudy (anonymous):

Okay... How much is -6+6?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Nope. It's NEGATIVE 6 plus POSITIVE 6 You can even swap it around: -6+6 is the same as +6-6 Which is more familiar as: 6-6 Which is... ? :-)

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

sorry I was doing something

OpenStudy (anonymous):

So am I. I'm trying to help you with a problem you posted.

OpenStudy (anonymous):

So what is -6+6?

OpenStudy (anonymous):

it would be 0 right??

OpenStudy (anonymous):

CORRECT!!!!! :-) So armed with that information, how can we simply: 3x -6 +6 = 15+6

OpenStudy (anonymous):

*simplify And don't let this throw you: "0" is a perfectly good number which we can show or get rid of at our discretion.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok so it would be 3x -0 = 15+6 right

OpenStudy (anonymous):

CORRECT AGAIN!!! :-) Now simplify the right hand side

OpenStudy (anonymous):

ok it would be 21 right :D

OpenStudy (anonymous):

RIGHT!! So what we have so far is: 3x -0 = 21 Now since we can add 0 or subtract 0 without making any difference at all, then it follows that we can also dismiss any 0 values without penalty, right? So that would leave us with: 3x = 21 Now that we have "isolated the terms" as it's called, we can now solve for "x" And we do that by dividing both sides by 3 PS: You can't add or subtract 3 from/to 3x, but you ARE allowed to multiply and divide 3x by 3)

OpenStudy (anonymous):

its x=7 isn't it :)

OpenStudy (anonymous):

YAY!!!!! POP THE CHAMPAGNE!!!! That's your answer! You've just "solved for x"!! :-)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

So you're all set now?

OpenStudy (anonymous):

yup thank you so much you had helped me a lot :-)

OpenStudy (anonymous):

No problem!! I'm very happy I could help! And thank you for the medal, as well! :-)

OpenStudy (anonymous):

:D your welcome

OpenStudy (anonymous):

correct